# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v1

@@ III-A2 Knowledge-augmented Dataset
Interest in incorporating knowledge into text-to-SQL tasks has increased significantly in recent years. BIRD [33] employs domain experts to annotate each text-to-SQL sample with external knowledge, categorized into Numeric Reasoning Knowledge, Domain Knowledge, Synonym Knowledge, and Value Illustration. Similarly, Spider-DK [39] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset: SELECT Columns Mentioned by Omission, Simple Inference Required, Synonyms Substitution in Cell Value Word, One Non-Cell Value Word Generate a Condition, and Easy to Conflict with Other Domains. Both studies found that human-annotated knowledge significantly improves SQL generation accuracy for samples requiring external domain knowledge. Additionally, SQUALL [44] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
@@ III-A3 Context-dependent Dataset
SParC [43] and CoSQL [35] explore context-dependent SQL generation by constructing a conversational database querying system. Unlike traditional text-to-SQL datasets that only have a single (question, SQL) pair for one example, SParC decomposes the (question, SQL) examples in the Spider dataset into multiple (sub-question, SQL) pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity. CoSQL, in comparison, involves conversational interactions in natural language, simulating real-life scenarios to increase complexity and diversity. Additionally, Spider-SS&CG [38] splits the NL question in the Spider dataset into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL model’s generalization ability on out-of-distribution samples.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL models with attacked or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [41] removes explicit schema-related words from the NL questions, while Spider-SYN [40] replaces them with manually selected synonyms. ADVETA [37] introduces Adversarial Table Perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. These perturbations lead to significant drops in accuracy, as a text-to-SQL model with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-B1 Content Matching-based Metrics
SQL content matching metrics focus on comparing the generated SQL query with the ground truth SQL query based on their structural and syntactic similarities.
  • Component Matching (CM) [13] evaluates the model performance by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score. Each component is decomposed into sets of sub-components and compared for an exact match, accounting for SQL components without order constraints.
  • Exact Matching (EM) [13] measures the percentage of examples whose predicted SQL query is equivalent to the ground truth SQL query. A predicted SQL is considered correct only if all its components, as described in Component Matching, match exactly with those of the ground truth query.
@@ III-B2 Execution-based Metrics
  • Execution Accuracy (EX) [13] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the ground truth queries.
  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries, which are the predicted SQL queries whose executed results exactly match the ground truth query. Thus, VES evaluates both the efficiency and accuracy of predicted SQL queries. For an evaluation dataset with $N$ examples, VES can be computed by:
@@
    where $\hat{Y}_{n}$ and $\hat{V}_{n}$ are the predicted SQL query and its executed results. $Y_{n}$ and $V_{n}$ are those of the corresponding ground truth SQL query. $\mathbbm{1}(V_{n},\hat{V}_{n})$ is an indicator function, where:
@@
    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database. [33] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following introduction.
@@ IV Methods
The implementation of LLM-based applications mostly relies on in-context learning (prompt engineering) [79] and fine-tuning [80] since the powerful proprietary models and well-architected open-source models are being released in large quantities [81, 82, 83]. The specific methods can generally be divided into two paradigms, in-context learning (ICL) and fine-tuning (FT). In this survey, we primarily focus on these text-to-SQL paradigms and will discuss them accordingly.
@@ IV-A In-context Learning
Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [28], also impacting the text-to-SQL task in different prompt styles [8, 46]. Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement. The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ is formulated as:
@@
where $\mathcal{Q}$ represents the user question. $\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$, where $\mathcal{C}=\{c_{1},c_{2},$...\} and tables $\mathcal{T}=\{t_{1},t_{2},$...\} represent the collection of various columns and tables, $\mathcal{K}$ is the potentially external knowledge (e.g. foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]). $\mathcal{I}$ represents the instruction for the text-to-SQL task, which performs indicative guidance to guide the LLMs for generating an accurate SQL query. $f(\cdot\mid\theta)$ is a LLM with parameter $\theta$. In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for implementation. Various well-designed methods have been adopted in the ICL paradigm for text-to-SQL. We group them into five categories $\mathbf{C}_{0:4}$, including: $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the details are shown in Tab. II.
@@
Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [84, 80, 85], which is widely recognized and achieves promising results. In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts. As introduce above, the formulated process of LLM-based text-to-SQL Eq. 3 can also represent zero-shot prompting, the input $\mathcal{P}_{0}$ can be obtained by concatenate $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
@@
To regulate the prompting process, we set the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate as the standard (trivial) prompt [30] for text-to-SQL.
TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL. The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Table III.
@@
TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time. The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Context Augmentation, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement. The method in multiple categories will be introduced respectively. *There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance. COT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [8], arXiv’24 [32]
@@
Zero-shot: Many research works [26, 27, 46]l utilize zero-shot prompting, focusing mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL. As an empirical evaluation, [26] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [86, 77, 87], which also presents the results for different prompt styles. The results indicate that prompt design is critical for performance, with error analysis, [26] propose more database content can harm the overall performance. Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation, [27] assesses its performance of text-to-SQL. With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL ability compared to the state-of-the-art (SOTA) models. For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
Primary key and foreign key carry contiguous knowledge of different tables, [49] study their impact by adding these keys on different designed database prompt style on zero-shot prompting results. A benchmark evaluation [8] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key. Apart from the foreign key, this study also explores zero-shot prompting combined with “no explanation” rule and the rule implication “Let’s think step by step”. Empowered by the annotated knowledge of human experts, [33] follow the standard prompting and obtain improvement by combining the provided oracle knowledge. With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46]. For zero-shot prompting optimization, [46] raises a challenge for designing an effective prompt template for LLMs; the former prompt construction lacks structure uniformity, which makes it hard to find out a concrete element within a prompt constructing template influences the performance of LLMs. They address this challenge by investigating a more unified series of prompt templates warping with different prefixes, infixes, and postfixes.
Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 88]. The input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq. 3:

### v2

@@ III-A2 Knowledge-augmented Dataset
Interest in incorporating domain-specific knowledge into text-to-SQL tasks has increased significantly in recent years. BIRD [33] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration. Similarly, Spider-DK [39] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [13]: SELECT columns mentioned by omission, simple inference required, synonyms substitution in cell value word, one non-cell value word generate a condition, and easy to conflict with other domains. Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge. Additionally, SQUALL [44] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
@@ III-A3 Context-dependent Dataset
SParC [43] and CoSQL [35] explore context-dependent SQL generation by constructing a conversational database querying system. Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity. CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity. Additionally, Spider-SS&CG [38] splits the NL question in the Spider dataset [13] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
@@ III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [41] removes explicit schema-related words from the NL questions, while Spider-SYN [40] replaces them with manually selected synonyms. ADVETA [37] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@ III-B1 Content Matching-based Metrics
SQL content matching metrics focus on comparing the predicted SQL query with the ground truth SQL query based on their structural and syntactic similarities.
  • Component Matching (CM) [13] evaluates the performance of text-to-SQL system by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score. Each component is decomposed into sets of sub-components and compared for an exact match, accounting for SQL components without order constraints.
  • Exact Matching (EM) [13] measures the percentage of examples whose predicted SQL query is identical to the ground truth SQL query. A predicted SQL is considered correct only if all its components, as described in CM, match exactly with those of the ground truth query.
@@ III-B2 Execution-based Metrics
  • Execution Accuracy (EX) [13] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the results obtained by the ground truth query.
  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries. A valid SQL query is a predicted SQL query whose executed results exactly match the ground truth results. Specifically, VES evaluates both the efficiency and accuracy of predicted SQL queries. For a text dataset with $N$ examples, VES can be computed by:
@@
    where $\hat{Y}_{n}$ and $\hat{V}_{n}$ are the predicted SQL query and its executed results and $Y_{n}$ and $V_{n}$ are the ground truth SQL query and its corresponding executed results, respectively. $\mathbbm{1}(V_{n},\hat{V}_{n})$ is an indicator function, where:
@@
    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database. BIRD benchmark [33] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following analysis.
@@ IV Methods
The implementation of current LLM-based applications mostly relies on in-context learning (ICL) (prompt engineering) [86, 87, 88] and fine-tuning (FT) [89, 90] paradigms since the powerful proprietary models and well-architected open-source models are being released in large quantities [85, 45, 91, 92, 93, 94]. LLM-based text-to-SQL systems follow these paradigms for implementation. In this survey, we will discuss them accordingly.
@@ IV-A In-context Learning
Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [95, 28], also impacting the SQL generation in different prompt styles [9, 46]. Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement. The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ can be formulated as:
@@
where $\mathcal{Q}$ represents the user question. $\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$, where $\mathcal{C}=\{c_{1},c_{2},$...\} and tables $\mathcal{T}=\{t_{1},t_{2},$...\} represent the collection of various columns and tables, $\mathcal{K}$ is the potentially external knowledge (e.g. foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]). $\mathcal{I}$ represents the instruction for the text-to-SQL task, which provides indicative guidance to trigger the LLMs for generating an accurate SQL query. $f(\cdot\mid\theta)$ is a LLM with parameter $\theta$. In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for generating the predicted SQL query. Various well-designed methods within the ICL paradigm have been adopted for LLM-based text-to-SQL tasks. We group them into five categories $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the representative methods of each category are given in Tab. II.
@@
Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [96, 89, 97], which is widely recognized and used in real-world applications. In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts (vanilla prompt engineering). As introduced above, Eq. 3 formulated the process of LLM-based text-to-SQL, which can also represent zero-shot prompting. The overall input $\mathcal{P}_{0}$ can be obtained by concatenating $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
@@
To regulate the prompting process, the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate is set as the standard (trivial) prompt [30] for text-to-SQL.
TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL. The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Tab. III.
@@
TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time. The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement. The method in multiple categories will be introduced respectively. *There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance. CoT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [9], arXiv’24 [32].
@@
Zero-shot: Many research works [7, 27, 46] utilize zero-shot prompting, studying mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL. As an empirical evaluation, [7] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [98, 84, 99] and the performance for different prompting styles. The results indicate that prompt design is critical for the performance, with error analysis, [7] propose more database content can harm the overall accuracy. Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation [100], [27] assesses its performance of text-to-SQL. With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL performance compared to the state-of-the-art PLM-based systems. For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
Primary keys and foreign keys carry contiguous knowledge of different tables. [49] studies their impact by incorporating these keys into various prompt styles with different database content to analyze zero-shot prompting results. A benchmark evaluation [9] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key. Apart from the foreign key, this study also explores zero-shot prompting combined with the rule implication “no explanation” to collect concise outputs. Empowered by the annotated external knowledge of human experts, [33] follows the standard prompting and achieves improvement by incorporating the provided annotated oracle knowledge.
With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46]. For zero-shot prompting optimization, [46] raises a challenge for designing an effective prompt template for LLMs; the former prompt construction lacks structure uniformity, which makes it hard to find out a concrete element within a prompt constructing template influences the performance of LLMs. They address this challenge by investigating a more unified series of prompt templates warping with different prefixes, infixes, and postfixes.
Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 101]. The overall input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq. 3: